Two bike trails have a total length of 20 miles. If the product of the length of the trails is 91, find the length of each trail.
Answer by checkley77(12569)

X+Y=20 OR X=20-Y
XY=91
(20-Y)Y=91
20Y-Y^2-91=0
Y^2-20Y+91=0
(Y-13)(Y-7)=0
Y-13=0
Y=13 ANS.
Y-7=0
Y=7 ANS.
PROOF:
7*13=91
91=91